# Dynamical Similarity in Multisymplectic Field Theory

### v3

@@ III.2 Hamiltonian Formalism
Recall that the Hamiltonian formalism of multisymplectic dynamics was developed on the restricted multimomentum bundle $J^{1}E^{*}$; in order to provide an analogous description for dissipative field theories, we shall work on the space
$\mathcal{S}^{*}:=J^{1}E^{*}\times_{M}\wedge^{m-1}T^{*}M\cong J^{1}E^{*}\times\mathbb{R}^{m}$
Introducing the projection $\bar{\tau}:\mathcal{S}^{*}\rightarrow M$, we obtain a bundle over $M$, of dimension $2m+n+nm$, upon which local coordinates are taken to be $(x^{\mu},y^{a},p^{\mu}_{a},s^{\mu})$.
Given a (hyper)regular Lagrangian system $(\mathcal{S},\Theta_{L})$, there exists a map $\mathcal{FL}:\mathcal{S}\rightarrow\mathcal{S}^{*}$, which acts on local coordinates according to
@@
This Legendre map then allows us to define the Hamiltonian $m$-form on $\mathcal{S}^{*}$ as $\Theta_{L}=\mathcal{FL}^{*}\Theta_{H}$; in local coordinates, $\Theta_{H}$ is given by
$\Theta_{H}=\left(ds^{\mu}-p^{\mu}_{a}\,dy^{a}\right)\,\wedge\,d^{m-1}x_{\mu}+H\,d^{m}x$ (32)
in which we have introduced the Hamiltonian function $H:\mathcal{S}^{*}\rightarrow\mathbb{R}$, related to the Lagrangian energy function according to $E_{L}=\mathcal{FL}^{*}H$. When studying the reduction of multisymplectic field theories to their frictional multicontact counterparts, it will be of benefit to work with the following 1-forms
@@
which is then used to construct the $(m+1)$-form $\Omega_{H}=\bar{d}\Theta_{H}$, as in (27).
A geometric formulation of the field equations may be given in terms of holonomic sections $\psi:M\rightarrow\mathcal{S}^{*}$, or locally decomposable, $\bar{\tau}$-transverse multivector fields $\boldsymbol{X}_{H}\in\mathfrak{X}^{m}(\mathcal{S}^{*})$. In the former case, we have
$\psi^{*}\Theta_{H}=0\quad\quad\quad\textrm{and}\quad\quad\quad\psi^{*}\iota_{Z}\,\Omega_{H}=0\quad\textrm{for all}\;Z\in\mathfrak{X}^{\infty}(\mathcal{S}^{*})$ (35)
whilst the corresponding equations for multivector fields are
$\iota_{\scriptscriptstyle\boldsymbol{X}_{H}}\Theta_{H}=0\quad\quad\quad\textrm{and}\quad\quad\quad\iota_{\scriptscriptstyle\boldsymbol{X}_{H}}\Omega_{H}=0$ (36)
If we decompose the multivector field $\boldsymbol{X}_{H}$ locally as
$\boldsymbol{X}_{H}=\bigwedge\limits_{\mu=1}^{m}\,\left(\frac{\partial}{\partial x^{\mu}}+A^{a}_{\mu}\,\frac{\partial}{\partial y^{a}}+B^{\nu}_{a\mu}\,\frac{\partial}{\partial p^{\nu}_{a}}+C^{\nu}_{\mu}\frac{\partial}{\partial s^{\nu}}\right)$ (37)
then the coordinate-free expressions (36) become
$A_{\mu}^{a}=\frac{\partial H}{\partial p^{\mu}_{a}}\quad\quad\quad B^{\mu}_{a\mu}=-\,\left(\frac{\partial H}{\partial y^{a}}+p^{\mu}_{a}\frac{\partial H}{\partial s^{\mu}}\right)\quad\quad\quad C^{\mu}_{\mu}=p^{\mu}_{a}\frac{\partial H}{\partial p^{\mu}_{a}}-H$ (38)
Let us suppose that $\psi:M\rightarrow\mathcal{S}^{*}$ is an integral section of $\boldsymbol{X}_{H}$; writing $\psi(x)=\left(x^{\mu},y^{a}(x),p^{\mu}_{a}(x),s^{\mu}(x)\right)$, the equations (35) are given locally by
@@
We shall refer to (38) and (39) as the Herglotz De-Donder Weyl equations for multivector fields and holonomic sections respectively; they are the appropriate modification of the standard Hamilton De-Donder Weyl equations, adapted to describe field theories which are frictional in nature [70]. Note that the last of the expressions in (39) is simply the condition $\partial_{\mu}s^{\mu}=L^{H}$, expressed in canonically conjugate variables.
III.3 The Scalar Field Revisited
In order to conclude our study of multisymplectic and multicontact geometry, we shall revisit the example of section (II.4), in which we considered a single free scalar field in a flat background spacetime. In an attempt to present an example which highlights the features of the preceding sections, but does not introduce further unnecessary complications, we shall content ourselves with a Lagrangian function which differs from that of (20) only by a simple linear term
$L(x^{\mu},y,y_{\mu},s^{\mu})=\frac{1}{2}\eta^{\mu\nu}y_{\mu}y_{\nu}-\frac{1}{2}m^{2}y^{2}+\lambda_{\mu}s^{\mu}$ (40)
in which $\lambda$ is some constant spacetime vector.
The Lagrangian 4-form $\Theta_{L}$ contains an additional contribution from the action density, and so differs somewhat from (22), that we calculated for the free field on $J^{1}E$
$\Theta_{L}=\left(ds^{\mu}-\eta^{\mu\nu}y_{\nu}\,dy\right)\,\wedge\,d^{3}x_{\mu}\,+\,\biggr(\,\underbrace{\frac{1}{2}\eta^{\mu\nu}y_{\mu}y_{\nu}+\frac{1}{2}m^{2}y^{2}-\lambda_{\mu}s^{\mu}}_{=\;E_{L}}\,\biggr)\,d^{4}x$ (41)
The Herglotz-Lagrange equations of motion for a holonomic section $\Psi:M\rightarrow\mathcal{S}=J^{1}E\times\mathbb{R}^{4}$ are given by
$\eta^{\mu\nu}\partial_{\mu}\partial_{\nu}y-m^{2}y=\eta^{\mu\nu}(\partial_{\mu}y)\lambda_{\nu}$
together with the requirement that $\partial_{\mu}s^{\mu}=L$; due to the presence of the additional term on the right-hand side of the above equation, the Lagrangian (40) describes a scalar field subject to a constant damping, whose strength is determined by $\lambda_{\mu}$.
It is a simple exercise to show that the Hamiltonian function $H:\mathcal{S}^{*}=J^{1}E^{*}\times\mathbb{R}^{4}\rightarrow\mathbb{R}$ is given by
$H(x^{\mu},y,p^{\mu},s^{\mu})=\frac{1}{2}\eta_{\mu\nu}p^{\mu}p^{\nu}+\frac{1}{2}m^{2}y^{2}-\lambda_{\mu}s^{\mu}$ (42)
In order to calculate $\Omega_{H}\in\Omega^{5}(\mathcal{S}^{*})$, we require an expression for the dissipation form $\sigma_{\Theta_{H}}$, which, from (34), we deduce to be
$\sigma_{\Theta_{H}}=\frac{\partial H}{\partial s^{\mu}}\,dx^{\mu}=-\lambda_{\mu}\,dx^{\mu}$ (43)
It thus follows that
$\Omega_{H}=d\Theta_{H}+\sigma_{\Theta_{H}}\wedge\,\Theta_{H}=dH\wedge d^{4}x-dp^{\mu}\wedge dy\wedge d^{3}x_{\mu}-\lambda_{\nu}dx^{\nu}\wedge(ds^{\mu}-p^{\mu}dy)\wedge d^{3}x_{\mu}$
Finally, suppose that
$\boldsymbol{X}_{H}=\bigwedge\limits_{\mu=0}^{3}\,\left(\frac{\partial}{\partial x^{\mu}}+A_{\mu}\,\frac{\partial}{\partial y}+B^{\nu}_{\mu}\,\frac{\partial}{\partial p^{\nu}}+C^{\nu}_{\mu}\frac{\partial}{\partial s^{\nu}}\right)$
is a locally decomposable $\bar{\tau}$-transverse multivector field satisfying $\iota_{\scriptscriptstyle\boldsymbol{X}_{H}}\Theta_{H}=0$ and $\iota_{\scriptscriptstyle\boldsymbol{X}_{H}}\Omega_{H}=0$; in this case, the component functions $A_{\mu}$, $B^{\nu}_{\mu}$, and $C^{\nu}_{\mu}$ satisfy
$A_{\mu}=\frac{\partial H}{\partial p^{\mu}}=\eta_{\mu\nu}p^{\nu}\quad\quad B^{\mu}_{\mu}=-\,\left(\frac{\partial H}{\partial y}+p^{\mu}\frac{\partial H}{\partial s^{\mu}}\right)=-m^{2}y+\lambda_{\mu}p^{\mu}\quad\quad C^{\mu}_{\mu}=p^{\mu}\frac{\partial H}{\partial p^{\mu}}-H=\frac{1}{2}\eta_{\mu\nu}p^{\mu}p^{\nu}-\frac{1}{2}m^{2}y^{2}+\lambda_{\mu}s^{\mu}$
@@ IV An Introduction to Dynamical Similarity
In this section, we shall present a brief introduction to the concept of and motivation behind dynamical similarity; in particular, we begin by illustrating how this specific class of symmetry arises within the familiar setting of symplectic geometry, and how the presence of a dynamical similarity points to a redundancy within our mathematical framework, which may ultimately be eliminated. The field-theoretic generalisation of this process of ‘contact reduction’ is the main objective of this work, and is presented in sections (IV.1) and (IV.2).
Consider a general symplectic manifold $(N,\omega)$, with $\textrm{dim}\,N=2n$; suppose also that $X_{H}\in\mathfrak{X}^{\infty}(N)$ is the (unique) Hamiltonian vector field associated with the function $H:N\rightarrow\mathbb{R}$ via
@@
for some $\Lambda\in\mathbb{R}$, which we refer to as the degree of the scaling symmetry. The first of these properties indicates that scaling symmetries belong to the set of non-strictly canonical transformations of $(N,\omega)$ [71]. Consider now the Lie bracket of $D$ and $X_{H}$, in particular
@@
Note that the vector field $D$ is not a symmetry of $X_{H}$ in the conventional sense i.e satisfying $D,X_{H}]=0$; instead, it acts to rescale the Hamiltonian vector field by an overall factor, which justifies our use of the term ‘scaling symmetry’.
A non-trivial corollary of this is that there exists an algebra of invariants of the scaling symmetry, whose dynamical evolution is both autonomous and insensitive to its action [24, 21]. In this sense, we may simplistically speak of a scaling symmetry as ‘mapping a physical solution into a second, indistinguishable one’. Having identified a scaling symmetry $D$, it is then possible to construct a space formed from the quotient of $N$ by the (one-dimensional) orbits of $D$; if we declare two points on the same $D$-orbit to be equivalent [^3] Note that this identification is the mathematical realisation of our qualitative statement regarding mapping equivalent physical solutions into one another., and denote the resulting quotient space $N/\sim$, then this space is of dimension 2n12n-1. Further, given a set of assumptions on the action of $D$, which are not overly restrictive (i.e that the one-parameter family of transformations generating $D$ acts freely and properly on $N$) $N/\sim$ is in fact a smooth manifold, with submersion $\pi:N\rightarrow N/\sim$, that admits a contact distribution $\xi:=\pi_{*}\,\textrm{ker}(\iota_{D}\omega)$ [72].
Within the Hamiltonian framework of symplectic mechanics, we typically work on the cotangent bundle $T^{*}Q$ over the $n$-dimensional configuration space $Q$ [73]; we then identify a scaling function $x:T^{*}Q\rightarrow\mathbb{R}$, which is simply a phase space function satisfying $Dx=x$. For many simple examples, $x$ can be chosen to be a single configuration space variable, with the result that the reduced space $T^{*}Q/\sim$ is isomorphic to $T^{*}\widetilde{Q}\times\mathbb{R}$, in which $\widetilde{Q}$ is a configuration space consisting of n1n-1 variables, with $x$ having been excised. Note that although $\textrm{dim}\,T^{*}\widetilde{Q}=2n-2$, we have eliminated only one degree of freedom, as the factor of $\mathbb{R}$ in $T^{*}\widetilde{Q}\times\mathbb{R}$ is spanned by the action $S$, which has now assumed the role of a dynamical variable.
Dynamical similarities are most prevalently encountered in theories in which our mathematical description makes reference (often in the name of calculational simplicity) to a variable associated with the global scale of the system [21]. The dynamical evolution of the set of observables is not affected by this scale parameter, and so may freely be removed from our ontology [74]. The factorisation of the reduced phase space [^4] This is, of course, not the most general final topology of the reduced space; however, it is both illustrative, and sufficiently common so as to warrant particular attention. into $T^{*}\widetilde{Q}\times\mathbb{R}$ gives a clear and intuitive interpretation, which we shall carry forward, when generalising these ideas to a field-theoretic discussion. A theory possessing a scaling symmetry is dynamically-equivalent to a second, simpler theory, consisting of a scale-invariant piece, together with an action-dependent component, which acts in a way that closely resembles friction [62].
The above statement contains two subtleties, which are of particular note. Firstly, the qualifying phrase ‘dynamically-equivalent’ is vital: the reduced theory faithfully reproduces the on-shell, observable dynamics of the original symplectic system; elsewhere, these two theories need not (and most likely will not) coincide. Additionally, we have described the action-dependent theory as exhibiting ‘friction-like’ effects. Typically, when describing a non-conservative system, one imagines that the mechanical energy lost is dissipated into other, empirically-inaccessible degrees of freedom [21, 75]. For a contact-reduced system, by contrast, the changes in energy may either be losses or gains; however, crucially, they arise as a result of the inaccessibility of the global scale variable to an intrinsic observer. For this reason, we employ the phrase ‘friction-like’, so as to ensure a clear distinction between our notion of non-conservative systems, and that of genuinely dissipative phenomena.
IV.1 Contact Reduction of Lagrangian Systems
Having presented a heuristic outline of the concept of contact reduction for theories of particles, we proceed with a detailed description of the analogous procedure within a field-theoretic framework, starting with the Lagrangian formalism. We will then discuss the geometry and topology of the reduced space, and show, through a series of simple examples, how the reduced theory has a particularly elegant physical interpretation, with potential implications for the quantisation procedure.
Let us consider a (regular) Lagrangian system $(J^{1}E,\Theta_{L})$ over the $d$-dimensional spacetime manifold $M$. Suppose, without loss of generality, that the Lagrangian function $L:J^{1}E\rightarrow\mathbb{R}$ depends upon the two field variables $\phi$ and $\psi$, and that, in order to preserve Poincaré invariance, the only dependence upon the spacetime coordinates $x^{\mu}$ is through that of $\phi(x)$ and $\psi(x)$.
We now identify a transformation $(x^{\mu},\phi,\psi)\mapsto(\lambda^{A}x^{\mu},\lambda^{B}\phi,\psi)$ on the covariant configuration space, under which the Lagrangian rescales as $L\mapsto\lambda^{\Lambda}L$ for $\Lambda\in\mathbb{R}$. This transformation may be implemented via the vector field
$\bar{\Sigma}=Ax^{\mu}\frac{\partial}{\partial x^{\mu}}+B\phi\frac{\partial}{\partial\phi}$ (45)
However, note that $\bar{\Sigma}\in\mathfrak{X}^{\infty}(E)$ is a vector field over $E$, and is therefore not directly applicable to the Lagrangian function. Consequently, we look to reproduce the effects of $\bar{\Sigma}$ by lifting it to a vector field $\Sigma$ on the first jet bundle [76]. It is relatively easy to see that
$\Sigma=B\phi\frac{\partial}{\partial\phi}+(B-A)\phi_{\mu}\frac{\partial}{\partial\phi_{\mu}}-A\psi_{\mu}\frac{\partial}{\partial\psi_{\mu}}$ (46)
correctly reproduces the transformation at the level of the jet bundle. Note that we have assumed that a single field $\phi$ undergoes a rescaling, whilst $\psi$ remains invariant. This may seem somewhat restrictive; however, it is always possible to perform a suitable change of coordinates, such that only one of the fields belonging to $E$ is rescaled; indeed, we shall see an example of this below.
Having deduced the vector field (46), we must then verify that $\Sigma$ satisfies
@@ IV.1 Contact Reduction of Lagrangian Systems
where, we recall, the $\theta_{L}^{\mu}$ are obtained from the Poincaré-Cartan form $\Theta_{L}$ according to (3). From our discussion above, it should be relatively clear that if $\Sigma$ is a scaling symmetry, then so too is any non-zero multiple of $\Sigma$. Consequently, the notion of a scaling symmetry is defined only up to an overall normalisation, and our choice (47) fixes this ambiguity.
In order to progress further with the reduction process, consider a Herglotz Lagrangian $L^{H}$, embedded within a multisymplectic manifold, of one dimension higher. We shall suppose that the Herglotz Lagrangian depends upon the $k$ scalar fields $\phi^{a}$ ($a=1,\,\cdots,k$), and possibly their first derivatives $\phi^{a}_{\mu}$, together with the action density $s^{\mu}$. Finally, we introduce a new field variable $\rho$, defined in such a way that
$\rho_{\mu}=-\,\frac{\partial L^{H}}{\partial s^{\mu}}$ (48)
Writing our multisymplectic Lagrangian $L$ in the following form
$L=e^{\rho}(L^{H}+\rho_{\mu}s^{\mu})$ (49)
a short calculation reveals that, when restricted to the reduced space, the equations of motion of $L^{H}$ and those of $L$ describe the same dynamics for all observable degrees of freedom.
@@
In light of this, suppose that we have identified a scaling symmetry $\Sigma$ of degree $\Lambda$, as outlined above. Referring to the decomposition (46), we then introduce the scaling function $\xi=\phi^{1/B}$, and, with the coordinate transformation $dy^{\mu}=\xi^{-A}\,dx^{\mu}$, the vector field $\Sigma$ is rendered of the form
$\Sigma=\xi\frac{\partial}{\partial\xi}+\xi_{\bar{\mu}}\frac{\partial}{\partial\xi_{\bar{\mu}}}$ (50)
@@
in which $\xi_{\bar{\mu}}$ denotes the velocity coordinate of $\xi$, taken with respect to $y$, that is $\xi_{\bar{\mu}}=\partial\xi/\partial y^{\mu}$. Finally, making the identification $\xi=e^{\rho}$, we find that the Lagrangian adopts the following form [63]
$L(\rho,\rho_{\bar{\mu}},\psi,\psi_{\bar{\mu}})=e^{\Lambda\rho}f(\rho_{\bar{\mu}},\psi,\psi_{\bar{\mu}})$ (51)
@@
for some function $f$ which, crucially, does not depend upon $\rho$. In these coordinates, the scaling symmetry is simply $\Sigma=\partial_{\rho}$.
In section (II.2), we presented a variational principle used to derive the Euler-Lagrange field equations; this method utilised the Poincaré-Cartan form $\Theta_{L}$, which is somewhat removed from the conventional construction of a variational principle, in which we carry out an integral of the Lagrangian. The origin of this apparent disparity lies in a formal feature of the multisymplectic variational principle, which is rarely discussed within the standard symplectic framework. The variational principle for multisymplectic systems makes use of objects known as Lepagean equivalents [77, 78]. In particular, the equations of motion are, in principle, deduced from the extremisation of the action
$\int_{M}\mathcal{L}\circ j^{1}\phi$
in which $\mathcal{L}\in\Omega^{d}(J^{1}E)$ is the Lagrangian density introduced in (1). However, this is a constrained calculation, since we are obliged to consider only those variations which lie within the space of holonomic sections [49]. It is convenient, therefore, to construct an equivalent variational principle, which allows the full space of sections to be considered without constraints. This is achieved through use of the form $\Theta_{L}$, which is a Lepagean equivalent of the Lagrangian density [79, 80].
Our interest in this seemingly inconsequential mathematical formality lies in the fact that, even though we use the Lepagean equivalent formulation for practical calculations, the equations of motion are, fundamentally, derived from an integral of the Lagrangian density. Consequently, should we wish map physical solutions into other, indistinguishable solutions, we must consider $\mathcal{L}$, and not just the Lagrangian function. This is of particular importance, given that we have made a reparameterisation of the spacetime coordinates, which has a direct effect on the volume form $\omega$. In particular, for a $d$-dimensional spacetime $M$, the redefinition $dy^{\mu}=\xi^{-A}\,dx^{\mu}$ leads to a corresponding change of $\omega\mapsto\xi^{A\cdot d}\,\omega$. We find that the relationship between the coordinate scaling $A$, the degree of the symmetry $\Lambda$, and $d$ is $\Lambda+A\cdot d=1$; this is a generalisation of the result of the time-reparameterisation of symplectic systems [72, 63], and the factor of the spacetime dimension enters the expression due to our reparameterising each of the $d$ coordinates equally.
The conclusion to be drawn from this discussion is that, in order to faithfully map between indistinguishable physical solutions, we must work with the $d$-form $L\,d^{d}x$, and not simply the Lagrangian function in isolation. From (51), we see that
$L\,d^{d}x\;\longrightarrow\;e^{\rho}f(\rho_{\bar{\mu}},\psi,\psi_{\bar{\mu}})\,d^{d}y:=\widehat{L}\,d^{d}y$ (52)
The Lagrangian function $\widehat{L}$ is then of a form reminiscent of (49); indeed, let us consider the Euler-Lagrange field equations (8) for $\rho$, applied to a Lagrangian of the form $\widehat{L}=e^{\rho}f(\rho_{\bar{\mu}},\psi,\psi_{\bar{\mu}})$; for ease of notation, we shall revert to $x^{\mu}$ to label spacetime coordinates, and refrain from placing bars over derivative indices.
@@
Comparing this expression for $f$ to the Lagrangian $L=e^{\rho}(L^{H}+\rho_{\mu}s^{\mu})$, and recalling that we require $\partial_{\mu}s^{\mu}=L^{H}$, it is clear that we should identify the action density and Herglotz Lagrangian as follows [^5] Strictly speaking, $\widehat{L}$ and $L^{H}$ belong to different spaces - the former being a function on $J^{1}E$, and the latter a function on $J^{1}E^{\prime}\times\mathbb{R}^{d}$ - and this should be reflected through appropriate use of submersions between the total space $E$ and the symmetry-reduced space $E^{\prime}$; however, for the sake of clarity, we have chosen to omit such maps.

### v4

@@ III.2 Hamiltonian Formalism
In parallel to the construction of $\mathcal{S}$ in (11), we introduce
$\mathcal{S}^{*}:=J^{1}E^{*}\times_{M}\wedge^{d-1}T^{*}M\cong J^{1}E^{*}\times\mathbb{R}^{d}$ (18)
with projection $\bar{\tau}:\mathcal{S}^{*}\rightarrow M$. Local coordinates on $\mathcal{S}^{*}$ are denoted $(x^{\mu},y^{a},p^{\mu}_{a},s^{\mu})$, upon which the Legendre map $\mathcal{FL}:\mathcal{S}\rightarrow\mathcal{S}^{*}$ acts according to
@@
Similarly, the Hamiltonian $d$-form on $\mathcal{S}^{*}$ is $\Theta_{L}=\mathcal{FL}^{*}\Theta_{H}$, and in local coordinates, we have
$\Theta_{H}=\left(ds^{\mu}-p^{\mu}_{a}\,dy^{a}\right)\,\wedge\,d^{d-1}x_{\mu}+H\,d^{d}x$ (20)
in which the Hamiltonian $H:\mathcal{S}^{*}\rightarrow\mathbb{R}$ is related to the Lagrangian energy function via $E_{L}=\mathcal{FL}^{*}H$. When studying the reduction of multisymplectic field theories to their frictional multicontact counterparts, it will be of benefit to work with the following 1-forms
@@
which is then used to construct the $(d+1)$-form $\Omega_{H}:=d\Theta_{H}+\sigma_{\Theta_{H}}\wedge\,\Theta_{H}$, as in (15). Given a holonomic section $\psi:M\rightarrow\mathcal{S}^{*}$, expressed as $\psi(x)=\left(x^{\mu},y^{a}(x),p^{\mu}_{a}(x),s^{\mu}(x)\right)$, the suitable modification of the Hamiltonian equations (10), adapted to action-dependent theories reads
@@
The final expression is simply the condition $\partial_{\mu}s^{\mu}=L^{H}$, expressed in canonically conjugate variables.
@@ IV An Introduction to Dynamical Similarity
We begin by motivating the concept of scaling symmetries for simple theories of particles. In the introduction, we provided a qualitative account of how many theories are described mathematically, in a way that represents not only the evolution of observable quantities, but contains additional redundant structure, often present to reduce calculational complexity. This can be made more precise by considering the phase space of a mechanical system, where the redundancy manifests itself as a vector field, generating scale transformations that leave all dimensionless observables invariant, while acting non-trivially on the canonical coordinates. Such symmetries are not, therefore, simple reparameterisations, which are merely alternative descriptions of the same underlying mathematical object. A scaling symmetry maps between points that are distinct elements of the system’s phase space. Such points are identified solely because all relational observables evolve in precisely the same manner.
To continue our example from the introduction, we consider a flat, homogeneous, isotropic Friedmann model, sourced by minimally-coupled scalar fields. The phase space of this system consists of the scale factor, the Hubble rate, the scalar field values and their momenta. However, the physical observables depend only on the Hubble rate, the scalar fields, and their velocities. The dynamical similarity within this system connects distinct points on the phase space manifold which have the same values of these observable parameters, and are thus physically indistinguishable. This is realised by a vector field, which rescales the phase space parameters in such a way so as to leave the empirically accessible ratios unchanged.
In order to provide a more mathematical description of these ideas, we consider a symplectic manifold $(N,\omega)$, with $\textrm{dim}\,N=2n$. Suppose that $X_{H}\in\mathfrak{X}^{\infty}(N)$ is the Hamiltonian vector field associated with the function $H:N\rightarrow\mathbb{R}$ via
@@
for some $\Lambda\in\mathbb{R}$, referred to as the degree of the scaling symmetry. The first of these properties indicates that scaling symmetries belong to the set of non-strictly canonical transformations of $(N,\omega)$ [42]. Together, these two conditions allow us to show that
@@
so that the vector field $D$ acts to rescale all phase space trajectories by the same non-zero factor. A non-trivial corollary of this is that there exists an algebra of invariants of the scaling symmetry, whose dynamical evolution is both autonomous and insensitive to its action [12, 17]. In this sense, a scaling symmetry maps one physical solution into a second, indistinguishable one.
Assuming that the flow of $D$ acts freely and properly on $N$, we have a well-defined quotient space $C:=N/\hskip-0.56905pt\sim$, in which two points are identified if they lie on the same $D$-orbit [43]. $C$ is a smooth manifold of dimension 2n12n-1, and the map $\pi:N\rightarrow C$, taking points of $N$ to their equivalence class under $D$-orbits is a submersion. Further, $\xi:=\pi_{*}\,\textrm{ker}(\iota_{D}\hskip 0.56905pt\omega)$ defines a contact distribution on $C$. Locally, we identify a scaling function $\rho:N\rightarrow\mathbb{R}$, defined to satisfy $D\rho=\rho\,$; the contact distribution $\xi$ is then represented as the kernel of the 1-form
$\pi^{*}\eta:=\frac{\iota_{D}\hskip 0.56905pt\omega}{\rho}$ (25)
In addition to the 1-form $\eta$, we may also introduce a contact Hamiltonian $H^{c}:C\rightarrow\mathbb{R}$ via
$\pi^{*}H^{c}:=\frac{H}{\rho^{\Lambda}}$ (26)
The system $(C,\eta,H^{c})$ is a regular contact Hamiltonian system, which faithfully reproduces the dynamics of the original symplectic system, without reference to $\rho$, which, physically, we identify with the inaccessible measure of global scale, whose presence is redundant, from the perspective of the evolution of the observables.
V Contact Reduction of Lagrangian Systems
Despite having introduced scaling symmetries within the Hamiltonian formalism, our field-theoretic discussion commences with the (regular) Lagrangian system $(J^{1}E,\Theta_{L})$ over the $d$-dimensional spacetime manifold $M$. In general, when a theory possesses a global measure of scale, whose presence is not required to describe the evolution of the observable degrees of freedom, there exist coordinates $(x^{\mu},\phi,\psi^{a})$ on $E$, with $a=1,\,\cdots,N$, such that under the transformation $(x^{\mu},\phi,\psi^{a})\mapsto(x^{\mu},\lambda^{K}\phi,\psi^{a})$, the Lagrangian function is rescaled according to $L\mapsto\lambda^{\Lambda}L$ for some $\Lambda\in\mathbb{R}$. In order to preserve Poincaré invariance, we have assumed that the only dependence of $L$ on the spacetime coordinates is through that of $\phi(x)$ and $\psi^{a}(x)$.
It is relatively clear that the vector field
$\Sigma:=K\left(\phi\frac{\partial}{\partial\phi}+\phi_{\mu}\frac{\partial}{\partial\phi_{\mu}}\right)$ (27)
reproduces the effect of the rescaling $(x^{\mu},\phi,\psi^{a})\mapsto(x^{\mu},\lambda^{K}\phi,\psi^{a})$; that is $\mathfrak{L}_{\Sigma}L=\Lambda L$. It must also be verified that $\Sigma$ satisfies
@@ V Contact Reduction of Lagrangian Systems
in which the $\theta_{L}^{\mu}$ were defined in (3). Note that the mathematical requirement for $\Sigma$ to constitute a scaling symmetry is simply that $\mathfrak{L}_{\Sigma}\,\theta_{L}^{\mu}=C\theta^{\mu}_{L}$, for some constant $C$. Imposing that $C=1$, as above, merely fixes an overall normalisation ambiguity.
In order to progress with the reduction process, consider a Herglotz Lagrangian $L^{H}$, embedded within a multisymplectic manifold, of one dimension higher, via the expression
$L(\rho,\rho_{\mu},\phi^{a},\phi^{a}_{\mu})=e^{\rho}(L^{H}(\phi^{a},\phi^{a}_{\mu},s^{\mu})+\rho_{\mu}s^{\mu})$ (29)
in which $\phi^{a}$ are a set of scalar fields, and we have made explicit all coordinate dependence. The variable $\rho$ is defined in such a way that
$\rho_{\mu}:=-\,\frac{\partial L^{H}}{\partial s^{\mu}}$ (30)
A short calculation then shows that the equations of motion of $L$ imply the corresponding Herglotz-Lagrange field equations for $L^{H}$, when $L$ is restricted to the subspace upon which $L^{H}$ is defined. Consequently, on this space, $L^{H}$ describes a dynamically equivalent theory to that of $L$, but does so without reference to $\rho$, which in this context, we interpret as the variable corresponding to an unobservable global scale.
@@
In light of this, upon identifying a scaling symmetry $\Sigma$ of degree $\Lambda$, as in (27), we define $\xi:=\phi^{1/K}$, satisfying $\Sigma\,\xi=\xi$. With this, the vector field $\Sigma$ is rendered of the form
$\Sigma=\xi\frac{\partial}{\partial\xi}+\xi_{\mu}\frac{\partial}{\partial\xi_{\mu}}$ (31)
@@
Making the identification $\xi=e^{\rho/\Lambda}$, we find that the Lagrangian adopts the following form [41]
$L(\rho,\rho_{\mu},\psi^{a},\psi^{a}_{\mu})=e^{\rho}f(\rho_{\mu},\psi^{a},\psi^{a}_{\mu})$ (32)
@@
for some function $f$ which, crucially, does not depend upon $\rho$. In these coordinates, the scaling symmetry is simply $\Sigma=\Lambda\hskip 0.56905pt\partial_{\rho}$. The Euler-Lagrange field equation for $\rho$ implies
@@
Comparing this expression for $f$ to the Lagrangian $L=e^{\rho}(L^{H}+\rho_{\mu}s^{\mu})$, and recalling that we require $\partial_{\mu}s^{\mu}=L^{H}$, suggests that we should identify the action density and Herglotz Lagrangian as follows